How many moles of Fe3O4 are required to supply enough iron to prepare 0.732 mol Fe2O3?
Answer #1

4 Fe3O4 + O2 --> 6 Fe2O3

Fe3O4/Fe2O3 = 4 : 6

= 2 : 3

= (2/3) : 1

Fe3O4 : Fe2O3 = (2/3) : 1

= 0.732 * [(2/3) : 1 ]

= (2/3)*0.732 : 0.732

= 0.488 : 0.732

Answer: 0.732 moles Fe2O3 are prepared from 0.488 moles Fe3O4